A marketing firm found that the equation S = 2.1A + 224 describes how the amount of sales S of a product depends on A, the amount spent on advertising the product. Both S and A are measured in thousands of dollars. For what amount, A, is the sales of a product at least 350 thousand dollars?


Answer by Theo(13342)

the formula is S = 2.1 * A + 224

when S = 350, the formula becomes 350 = 2.1 * A + 224

subtract 224 from both sides of this equation and then divide both sides of this equation by 2.1 to get:

(350 - 224) / 2.1 = A

solve for A to get A = 60.

so, if 60 thousand is spent on advertising, then the amount of sales can be expected to be 350,000.

your formula becomes S = 2.1 * 60 + 224 = 350.

all dollar figures are in thousands of dollars, including, presumably, the 224.